OpenStudy (anonymous):

Simplify the expression -5+i/2i Can you also show me the steps of how to do it

OpenStudy (anonymous):

\[\frac{ -5+i }{ 2i }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, extend fraction by i. use this rule: i*i=-1 (definition) -5i -1 ------ 2*-1

OpenStudy (anonymous):

okay so -5(-1)-1 --------- -2 right? or is that an answer that you wrote

OpenStudy (anonymous):

original equation: -5 + i ----- 2 i extend numerator and denominator by i: i (-5 + i) --------- i (2 i) calculate: -5 i + i*i --------- 2 i * i use the rule i*i = -1: -5i +(-1) -------- 2 (-1)

OpenStudy (anonymous):

oh so the answer would be -5i-1 -------- -2

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

you can also multiply by -2 to simplify a bit...

OpenStudy (anonymous):

Yay thank you so much

OpenStudy (anonymous):

no issue

OpenStudy (anonymous):

5i + 1 ----- 2 should be the same thing

OpenStudy (anonymous):

yes it should :)